TIME-RESOLVED INFRARED DIODE LASER SPECTROSCOPY OF THE Co(CO)_{2} RADICAL

MOTOKI NAKASHIMA, SEIKI IKEDA, TOSHIHIDE HIKIDA, \underline{KEIICHI TANAKA} and TAKE-HIKO TANAKA, Department of Chemistry, Faculty of Sciences, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

Rovibrational transitions of the ν_{3} band (C-O anti-symmetric stretch) of the Co(CO)_{2} radical were observed in the frequency region of 1944 - 1964 cm^{−1}. The Co(CO)_{2} radical was produced by the 248 nm excimer laser photolysis and the transient absorption signal was detected by time-resolved infrared diode laser spectroscopy. More than 160 absorption lines were assigned to the ν_{3} fundamental band. The ν_{3} band consists of two series of spectral lines spaced approximately by four times the rotational constant, due to the spin statistics of identical C and O nuclei, and the electronic ground state was confirmed to be X^{2}Δg. Molecular constants including the band origin ν_{0}, rotational constant, and centrifugal distortion constant were derived from the observed transitions. Hot bands originating from the ν_{7} and 2ν_{7} vibrational excited states, where ν_{7} refers to the C-Co-C bending mode, were also observed. Because of the rovibronic interaction, each of the observed lines for the ν_{3} + ν_{7} ← ν_{7} and ν_{3} + 2ν_{7} ← 2ν_{7} hot bands was split into two and three components, respectively. Ab initio calculation in BP86/6-311+G* level suggests that the Co(CO)_{2} radical is slightly bent and has the X^{2}A_{1} electronic ground state. The present experiment shows that the Co(CO)_{2} radical is linear in average, and it may probably be a quasi-linear molecule in which the ground state has energy higher than the top of the potential bump.